4. A sprinter nuns at constant speed 9 ms around a circular track with a centripetal asselsration of magnitude 26m (a) What is the track s radius?(b) What is the magnitude of angular velocity of the sprinter? (c) What is the magnitude of the sprinters angular acceleration? (d) How long des it take to complete one revolution (this time is called period of motion)? (e) How many revolutions per minute does the sprinter make?
Answer #1

a] centripetal acceleration a = v^2/r

r = v^2/a = 9.2^2/7.6 = 11.14 m answer

b] angular velocity = v/r = 9.2/11.14 = 0.826 rad/s

c] angular acceleration = 0

d] Period = 2pi*r/v = 2pi*11.14/9.2 = 7.61 s

e] revolution per minute = 60 /7.61 =7.88 rev/min